# RAPTOR: Practical Numerical Profiling of Scientific Applications

### v1

@@ 3.4. Runtime
The RAPTOR runtime executes floating-point operations in the specified precisions and collects data for analysis. This can be done using hardware, when the target precision is available on the CPU. Alternatively, for arbitrary precisions, we use GNU’s MPFR library (Fousse et al., 2007) for floating-point emulation, which provides feature rich and efficient implementation for arithmetic and mathematical functions. Figure 5 sketches the MPFR-based wrappers in RAPTOR.
@@ 3.5. Operation Modes
Figure 5. Implementation of the RAPTOR runtime.
In this mode, the floating-point values that cross the boundary between the RAPTOR runtime and the user code (i.e. parameters to the runtime calls and return values) are represented using the pre-truncated floating-point type. For example, in the case in Figure 4a, the arguments to the _raptor_add_f32 function and its return value will be valid floating-point numbers of type float. Figure 5a shows the runtime implementation. Each time we perform any floating-point operation on these values, they first need to be transformed to the desired truncated precision (mpfr_set), the operations performed in truncated precision, and then expanded again to the original floating-point type (mpfr_get) [^17] This approach ensures correctness in adjunct code regions, e.g., calls to printf.. Since the values are converted back to the original type after each operation, there is no way of tracking how values flow through the computation. So, this mode is useful for collecting statistics about errors in individual operations (hence, op-mode), and the number thereof.
@@
Figure 5b shows a simplified version of the mem-mode runtime implementation. We use the bits in the floating-point type to store an integer identifier, which is used to recover a struct (_raptor_fp) containing the MPFR variable among other (customizable) bookkeeping data. Here, we add a double-precision shadow variable to the struct and update it with full-precision operations alongside the truncated MPFR variable. Hence, we can monitor the deviation between truncated variables and the (FP64) shadow variables for every operation, allowing us to set/monitor error threshold (cf. Section 6.3) and correlate them back to source code locations.
@@
MPFR Optimization
[…]
@@ 4. Precision Considerations in Non-linear Multi-Physics Workloads
To evaluate RAPTOR in the upcoming sections, we require a set of realistic application use cases. Our intent is to select a class of applications and numerical methods that do not have mathematical foundations that could otherwise help us understand and predict behavior under mixed-precision. We use Flash-X (Dubey et al., 2022), a scientific software that is used in multiple science domains. Flash-X is particularly well suited for this study because its target applications are non-linear and coupled multi-physics problems which are sensitive to numerical perturbations. Additionally, Flash-X uses Adaptive Mesh Refinement (AMR), a computational technique used to dynamically adjust the resolution of a simulation mesh based on the evolving features of the solution. Instead of using a uniform grid everywhere, AMR refines the mesh in regions where fine detail is needed (e.g., near steep gradients or discontinuities such as shocks and interfaces) and coarsens it where the solution is smooth. This allows simulations to achieve high accuracy while reducing computational cost by concentrating resources where they matter most. In the context of this study, the use of AMR permits us to exercise scientific intuition for adjusting the precision because AMR relies heavily on user-defined or computed thresholds for the refinement.
(a) The radial shock in the Sedov blast simulation
(b) Shock along vertical plane is Sod shock-tube simulation
Figure 6. Compressible hydrodynamics workloads with varying AMR block sizes. Each block contains a fixed number of grid cells. The background color qualitatively represents pressure variations, with darker regions indicating higher pressure and lighter regions indicating lower pressure.
@@ 4.1. Experiment Design
Flash-X uses a version of AMR where the physical domain is divided into blocks that are spatially organized in an octree. The physical size of the blocks at a given level on the tree is identical; blocks a level above are twice the size along each dimension, while those at the level below are half the size. Figure 6(a) illustrates this for a 2D domain. The number of cells is identical in every block.
Our first set of experiments (see Section 6.1 and Section 6.2) is predicated on the intuition that the refinement criteria mentioned above ensure that all blocks that need the highest level of accuracy will be at the finest refinement level in the octree. The blocks at coarser levels in the tree will have smoother solutions implying that they may have a potential for adjusting the precision. In applications that simulate shock waves, the highest refinement will follow the progress of the shock, while in the applications which simulate multiphase flows, the highest resolution will follow the phase boundaries. We can explore this hypothesis by applying truncation on blocks that are at higher levels in the tree. We can investigate the impact of precision by varying the levels at which we commence the truncation, or by truncating specific physics solvers, and correlating the resulting error to our knowledge of the physical regimes being explored in each simulation. These experiments use RAPTOR’s op-mode with function and file-level scoping.
@@ 4.2. Flash-X Workloads and Hypotheses
Flash-X can be configured for different applications with different underlying physics and solvers. Here, we select applications in two different physical regimes, i.e., with compressible and incompressible hydrodynamics solvers. In workloads with compressible hydrodynamics, shocks (sharp, nearly discontinuous change in the properties of fluid over a very small region of space) develop due to discontinuities in the initial conditions. We select two applications with different shock profiles. The first case, Sedov blast wave (Sedov, 1993), is initialized with a pressure spike in the center of the domain where the shock moves out in the radial direction. The regions away from the shock are more or less quiescent (see Figure 6(a)). The second case is the Sod shock tube (Sod, 1978) where there is a jump in density along the vertical plane. The shock wave moves in one direction while the rarefaction wave moves in the opposite direction, see Figure 6(b). For these workloads, we theorize:
@@
Hypothesis 2: EOS, used in the simulation, is table-based, and is therefore the most likely candidate for reducing precision.
@@
This Bubble simulation is governed by key dimensionless parameters for fluid and flow configuration. With $\rho^{\prime}=1000$ and $\mu^{\prime}=100$, we denote the density and viscosity ratios between water and air, respectively. These two parameters control the scaling of the fluid forces between the two phases. Furthermore, the Reynolds number (Re), Froude number (Fr), and Weber number (We) characterize the relative importance of inertial, gravitational, and surface tension forces. These numbers are set for water phase and scaled in air by the density and viscosity ratios, and we use $\text{Fr}=1$ and $\text{We}=125$. This configuration corresponds to a benchmark case reported in (Dhruv, 2024, 2023).
@@ 5. Experimental Setup
Here, we briefly describe details about the sub-components of RAPTOR, the execution environment, and how to read the figures.
(a) Between the 1st and 2nd panels, the error drops by seven orders of magnitude for mantissas smaller than 18 bits and remains constant at $10^{-15}$ for mantissas larger than 18 bits. Restricting truncation to coarser levels in 3rd/4th panels does not significantly change the error plot. Gray lines shadow other panels’ error to ease the comparison.
@@
Figure 7. Truncating hydrodynamics in the Sedov blast wave simulation (left) and Sod shock tube simulation (right). The primary $y$-axis shows the $L1$ error norm of the fluid density compared to the full-precision simulation, as computed by Flash-X’ serial output comparison utility sfocu. Results are plotted with lines (not points) for clarity. Gray lines, showing the errors from the other panels, are included to emphasize small differences in error from one panel to the next. Black vertical lines mark half, float, and double precision mantissa sizes with 10, 23, and 52 bits, respectively. In these plots, each panel represents a different refinement cutoff level beyond which truncation was disabled (i.e., $M-0$: truncate everything; $M-1$: disable truncation for most refined AMR level; $M-2$: disable truncation for two most refined levels; etc.). The bars in the background of each plot show the number of truncated operations (red) and full-precision operations (orange) stacked on top. As the refinement cutoff level is coarsened, the portion of truncated operations shrinks as truncation is restricted to coarser levels. Section 7.2 demonstrates how the operation counts can be used with a simple performance model to predict speedups. For small mantissas towards the left of each panel, the operation counts tend to deviate from the mean. The counts are perturbed when the AMR algorithm refines the grid to compensate for inaccuracies caused by aggressive truncation.
5.1. Versioning and Execution Environment
@@
For LLVM, we use tag 20.1.0-init and enhance it with a patch to add function wrappers for Fortran. Our fork of Enzyme is based on the trunc-trace branch. We utilize Spack (Gamblin et al., 2015) to install MPFR in version 4.2.1 and to install a Bubble dependency, called Hypre (v2.31.0). We compile additional Flash-X dependencies (HDF5 v1.14.6, OpenMPI v5.0.6, AmReX v24.08, MA28 v1.0.0) needed for the experiments. To gain access to the compressible hydrodynamics applications, we fork Flash-X’ main branch, state of 2025/01/24202020Flash-X carries an Apache 2.0 license but is maintained privately; Please email flash-x@lists.cels.anl.gov with your Github ID to get access (reference: #ec2a750).. For the Bubble application, we fork Flash-X’ state of 2025/04/14212121For reference: #cf6c018..
@@
We run all experiments in the default configuration for Flash-X: CPU-only and MPI-only. The Sod, Sedov, and Cellular applications execute sufficiently fast, and hence we use a single MPI process. The Bubble workload is computationally more demanding, requiring us to execute it with 32 MPI ranks, while still confining it to a single compute node. However, the parallelization across ranks does not affect the outcome of our experiments [^22] No MPI collectives are called within truncated sections; the domain is split over MPI processes and the physics routines that we truncate operate locally in each cell.. Our cluster’s compute nodes are equipped with dual-socket AMD EPYC 7773X processors and 1024 GiB of DDR4 main memory.
@@
We employ the same methodology for each experiment: First, we evenly lower precision in a given module across the entire mesh to set the baseline of expected behavior. Then, we start with a very small mantissa and gradually increase its size getting an error estimate for each instance. Once the baseline is established we repeat the above process by lowering precision on blocks in the target physics solver at levels $M-1$, $M-2$, and $M-3$ respectively where $M$ is the maximum refinement level.
@@ 6. RAPTOR Case Studies: Experimental Results and Insights for the four AMR Workloads
Hereafter, we describe our results from analyzing the Flash-X workloads with RAPTOR. We explore three truncation modes: (1) Global truncation, where numerical values are truncated uniformly across the entire domain, (2) Selective truncation with AMR, where truncation is applied only on levels coarser than ${M-l}$, with $l$ controlling the cutoff depth in the AMR hierarchy, and (3) Selective truncation of a complete physics module.
@@ 6.1. Compressible hydrodynamics applications
Figures 7(a) and 7(b) show the results of the hydro experiments. As expected, the Sedov problem exhibits a very robust tolerance to lowered precision at $M-1$ as demonstrated in the second panel of Figure 7(a), which shows a much lower error for any size of mantissa compared to the error shown in the first panel. For mantissas with less than 28 bits, excluding the finest AMR blocks from truncation results in an error reduction of more than 7 orders-of-magnitude. For mantissas larger than 28 bits, the error remains constant and approaches the error measured for full truncation. This error floor appears as the simulation result drifts over time due to small inaccuracies in truncated floating-point operations. Excluding the second level of AMR blocks from truncation does not significantly improve the error of the simulation, as demonstrated by the unchanging error in Panel 3. Only when we exclude the finest three levels of AMR blocks from truncation do we see that the constant error floor drops by another half order of magnitude as shown in Panel 4.
@@
The colored bars in the background of the figure panels count the number of truncated and full-precision floating-point operations occurring for each simulation. In the fourth panel, the workload is comprised almost entirely of full-precision operations, with truncated operations making up less than $0.4\%$. In the third and second panel, the truncated operations make up progressively larger portions of the workload as the Hydro module is truncated for more blocks in the domain. Finally, in the first panel, the truncated operations make up more than 80% of the floating-point workload. Additionally, for mantissas smaller than 10 bits, the bars show an instability, with the overall flop count going up and down. This occurs because the low-precision operations are influencing the decisions of the AMR algorithm, which is responsible for dynamically refining the grid where needed. To be clear, it is not the algorithm itself which is working with truncated precision. Rather, the algorithm notices imprecise blocks and decides to refine them in order to maintain numerical accuracy and convergence in the simulation.
@@
Figure 7(b) shows the error plot for the Sod problem. Comparing the first two panels, we observe that excluding the finest AMR blocks from truncation results in an improvement in the overall error for mantissa sizes smaller than 18 bits. This improvement results in an almost order-of-magnitude difference in error between the first and second panel for simulations using 4 bits of mantissa. Panel 3 shows that excluding another level of AMR blocks from truncation results in an error improvement across the board, with the difference starting at about half an order-of-magnitude for large mantissas with 48 bits and increasing to almost an order-of-magnitude for small mantissas with 10 bits.
@@
Excluding refined blocks from truncation improves the error more for the Sedov problem than it does for the Sod problem, validating Hypothesis 1. For Sedov, running just the finest blocks at full precision results in an improvement of 7 orders-of-magnitude compared to truncating all blocks. For Sod, the same strategy decreases the error by an order-of-magnitude at most. The figure for Sod has one less panel, because eventually no leaf blocks remain at the $M-3$ level in the simulation and therefore there is no error.
@@
In Figure 7(b), an anomalous behavior is visible for mantissas bits between 4-6, where the errors are comparable to those obtained for mantissas with 20 bits. The reason for that anomaly is written in the logfile of the Flash-X run, i.e., it shows a higher number of leaf blocks (the blocks on which the solution evolves) which implies that many more blocks meet the refinement criterion at such low mantissa bits. Consequently, any advantage from lowering the precision is more than offset by the increase of refined blocks. It highlights the importance of a careful cost-benefit analysis of any performance-related tweaks in the application and showcases how RAPTOR can assist with such an analysis.
@@
With Cellular we turn our attention to the possibility of truncating an entire physics module. The EOS module is the obvious choice here because the ordinary differential equations (ODEs) in the Burn module are particularly stiff and sensitive to numerical perturbation. Hypothesis 2 is based on the assumption that the computation can work with reduced precision because it extrapolates from a table look-up. However, this hypothesis is proven to be incorrect because the Newton-Raphson algorithm, used to extrapolate from the table look-up values to actual conditions in the simulation, does not converge within the specified number of iterations when the mantissa is truncated to less than 42 bits. In an attempt to get around this problem, we decrease the tolerance for convergence and increase the permitted number of iterations. Yet, we fail to get convergence for any meaningful workload, providing us with a counterexample for our hypothesis. This experiment in particular highlights the usefulness of RAPTOR in numerical profiling because scientific intuition is not always reliable. Any direct modification of code to run such experiments would be time consuming and not as informative.

### v2

@@ 3.4. Runtime
The RAPTOR runtime executes floating-point operations in the specified precisions and collects data for analysis. This can be done using hardware, when the target precision is available on the CPU. Alternatively, for arbitrary precisions, we use the GNU MPFR library (Fousse et al., 2007) for floating-point emulation, which provides feature rich and efficient implementations of arithmetic and mathematical functions. Figure 5 sketches the MPFR-based wrappers in RAPTOR.
@@ 3.5. Operation Modes
In this mode, the floating-point values that cross the boundary between the RAPTOR runtime and the user code (i.e. parameters to the runtime calls and return values) are represented using the pre-truncated floating-point type. For example, in the case in Figure 4, the arguments to the _raptor_add_f32 function and its return value will be valid floating-point numbers of type float. Figure 5 shows the runtime implementation. Each time we perform any floating-point operation on these values, they first need to be transformed to the desired truncated precision (mpfr_set), the operations performed in truncated precision, and then expanded again to the original floating-point type (mpfr_get). [^13] This approach ensures correctness in adjunct code regions, e.g., calls to printf. Since the values are converted back to the original type after each operation, there is no way of tracking how values flow through the computation. So, this mode is useful for collecting statistics about errors in individual operations (hence, op-mode), and the number thereof.
[…]
@@
Figure 5 shows a simplified version of the mem-mode runtime implementation. We use the bits in the floating-point type to store an integer identifier, which is used to recover a struct (_raptor_fp) containing the MPFR variable among other (customizable) bookkeeping data. Here, we add a double-precision shadow variable to the struct and update it with full-precision operations alongside the truncated MPFR variable. Hence, we can monitor the deviation between truncated variables and the (FP64) shadow variables for every operation, allowing us to set/monitor error thresholds (cf. Section 6.3) and correlate them back to source code locations.
@@
(a) The radial shock in the Sedov blast simulation
(b) Shock along vertical plane in the Sod shock-tube simulation
Figure 6. Compressible hydrodynamics workloads with varying AMR block sizes. Each block contains a fixed number of mesh cells. The background color qualitatively represents pressure variations, with darker regions indicating higher pressure and lighter regions indicating lower pressure.:
3.6. Compatibility
RAPTORrecognizes OpenMP directives and correctly truncates operations within nested OpenMP parallel constructs within the same device.
In addition to truncating applications for CPU, RAPTOR can truncate code for GPUs (AMD and NVidia). We have successfully tested OpenMP offloading and CUDA code. Due to the lack of GPU support in arbitrary precision libraries (such as MPFR), truncation is currently only possible to native types available in the GPU hardware, such as float or half precision floating-point.
In general, RAPTOR’s op-mode and MPI do not interfere with one another and truncation continues to work for any application with one or more MPI ranks. Most MPI operations only involve message passing and therefore require no special handling. However, RAPTOR does not implicitly truncate MPI reductions or any other MPI functions that perform floating-point operations. If e.g. truncated MPI Allreduce is needed, a custom reduction operation can be implemented, which in turn can be truncated using RAPTOR. Mem-mode can only be used on shared-memory systems and without MPI reductions. [^16] In op-mode, reductions do not get truncated automatically. In mem-mode, reductions will crash the application when they change the pointers to the shadow storage. Otherwise, additional MPI communication and handling would be required to transfer the shadow storage and adjust the pointers.
@@ 4. Precision Considerations in Non-linear Multi-Physics Workloads
To evaluate RAPTOR in the upcoming sections, we require a set of realistic application use cases. Our intent is to select a class of applications and numerical methods that do not have mathematical foundations that could otherwise predict mixed-precision behavior. We use Flash-X (Dubey et al., 2022), a scientific software that is used in multiple science domains. Flash-X is well suited for this study because its target applications are non-linear and coupled multi-physics problems which are sensitive to numerical perturbations. Additionally, Flash-X uses Adaptive Mesh Refinement (AMR), a computational technique used to dynamically adjust the resolution of a simulation mesh based on the evolving features of the solution. Instead of using a uniform grid everywhere, AMR refines the mesh in regions where fine detail is needed (e.g., near steep gradients or discontinuities such as shocks and interfaces) and coarsens it where the solution is smooth. This allows simulations to achieve high accuracy while reducing computational cost by concentrating resources where they matter most. Here, the use of AMR permits us to exercise scientific intuition for adjusting the precision because AMR relies heavily on user-defined or computed thresholds for the refinement.
@@ 4.1. Experiment Design
Flash-Xuses a version of AMR where the physical domain is divided into blocks that are spatially organized in an octree. The physical size of the blocks at a given level on the tree is identical; blocks one level above are twice the size along each dimension, while those one level below are half the size. Figure 6(a) illustrates this for a 2D domain. The number of cells is identical in every block.
Our first set of experiments (see Section 6.1 and Section 6.2) is predicated on the intuition that the refinement criteria mentioned above ensure that all blocks that need the highest level of accuracy will be at the finest refinement level in the octree. The blocks at coarser levels in the tree will have smoother solutions implying that they may have a potential for adjusting the precision. In applications that simulate shock waves, the highest refinement will follow the progress of the shock, while in the applications which simulate multiphase flows, the highest resolution will follow the phase boundaries. We can investigate the impact of precision by varying the levels at which we commence the truncation, or by truncating specific physics solvers, and correlating the resulting error to our knowledge of the physical regimes being explored in each simulation. These experiments use RAPTOR’s op-mode with function and file-level scoping.
@@ 4.2. Flash-XWorkloads and Hypotheses
Flash-Xcan be configured for different applications with different underlying physics and solvers. Here, we select applications in two different physical regimes, i.e., with compressible and incompressible hydrodynamics solvers. In workloads with compressible hydrodynamics, shocks (sharp, nearly discontinuous change in the properties of fluid over a very small region of space) develop due to discontinuities in the initial conditions. We select two applications with different shock profiles. The first case, Sedov blast wave (Sedov, 2018), is initialized with a pressure spike in the center of the domain where the shock moves out in the radial direction. The regions away from the shock are more or less quiescent (see Figure 6(a)). The second case is the Sod shock tube (Sod, 1978) where there is a jump in density along the vertical plane. The shock wave moves in one direction while the rarefaction wave moves in the opposite direction, see Figure 6(b). For these workloads, we theorize:
@@
Hypothesis 2: The EOS, used in the simulation, is table-based and is therefore the most likely candidate for reducing precision.
@@
The Bubble simulation is governed by key dimensionless parameters for fluid and flow configuration. With $\rho^{\prime}=1000$ and $\mu^{\prime}=100$, we denote the density and viscosity ratios between water and air, respectively. These two parameters control the scaling of the fluid forces between the two phases. Furthermore, the Reynolds number (Re), Froude number (Fr), and Weber number (We) characterize the relative importance of inertial, gravitational, and surface tension forces. These numbers are set for water phase and scaled in air by the density and viscosity ratios, and we use $\text{Fr}=1$ and $\text{We}=125$. This configuration corresponds to a benchmark case reported in (Dhruv, 2024; Akash Dhruv, 2023).
@@ 5. Experimental Setup
Here, we briefly discuss versions and details of the sub-components of RAPTOR. We also describe the execution environment and lay out our strategy for running the experiments.
(a) Between the 1st and 2nd panels, the error drops by seven orders of magnitude for mantissas smaller than 18 bits and remains constant at $10^{-15}$ for mantissas larger than 18 bits. Restricting truncation to coarser levels in the 3rd and 4th panels does not significantly change the error plot.
@@
Figure 7. Truncating hydrodynamics in the Sedov blast wave simulation (left) and Sod shock tube simulation (right). The primary $y$-axis shows the $L1$ error norm of the fluid density compared to the full-precision simulation, as computed by Flash-X’ serial output comparison utility sfocu. Results are plotted with lines (not points) for clarity. Gray lines, showing the errors from the other panels, are included to emphasize small differences in error from one panel to the next. Black vertical lines mark half, float, and double precision mantissa sizes with 10, 23, and 52 bits, respectively. In these plots, each panel represents a different refinement cutoff level beyond which truncation was disabled (i.e., M0M-0: truncate everything; M1M-1: disable truncation for most refined AMR level; M2M-2: disable truncation for two most refined levels; etc.). The bars in the background of each plot show the number of truncated operations (red) and full-precision operations (orange) stacked on top. As the refinement cutoff level is coarsened, the portion of truncated operations shrinks as truncation is restricted to coarser levels. Section 7.2 demonstrates how the operation counts can be used with a simple performance model to predict speedups. For small mantissas towards the left of each panel, the operation counts tend to deviate from the mean. The counts are perturbed when the AMR algorithm refines the mesh to compensate for inaccuracies caused by aggressive truncation. :
@@
For LLVM, we use version 20, which offers a major improvement of the flang compiler over previous releases. Our fork of Enzyme is based on the trunc-trace branch. We utilize Spack (Gamblin et al., 2015) to install MPFR version 4.2.1 and to install a Bubble dependency, called Hypre (v2.31.0). We compile additional Flash-X dependencies (HDF5 v1.14.6, OpenMPI v5.0.6, AmReX v24.08, MA28 v1.0.0) needed for the experiments. To gain access to the compressible hydrodynamics applications, we fork Flash-X’ main branch, state of 2025/01/24.171717Flash-Xcarries an Apache 2.0 license but is maintained privately; Please email flash-x@lists.cels.anl.gov with your Github ID to get access (reference: #ec2a750). For the Bubble application, we fork Flash-X’ state of 2025/04/14.181818For reference: #cf6c018.
@@
We run all experiments in the default configuration for Flash-X: CPU-only and MPI-only. The Sod, Sedov, and Cellular applications execute sufficiently fast, and hence we use a single MPI process. The Bubble workload is computationally more demanding, requiring us to execute it with 32 MPI ranks, while still confining it to a single compute node. However, the parallelization across ranks does not affect the outcome of our experiments. [^19] No MPI collectives are called within truncated sections; the domain is split over MPI processes and the physics routines that we truncate operate locally in each cell. Our cluster’s compute nodes are equipped with dual-socket AMD EPYC 7773X processors and 1024 GiB of DDR4 main memory.
@@
We employ the same methodology for each experiment: First, we evenly lower precision in a given module across the entire mesh to set the baseline of expected behavior. Then, we start with a very small mantissa and gradually increase its size getting an error estimate for each instance. Once the baseline is established, we repeat the above process by lowering precision on blocks in the target physics solver at levels M1M-1, M2M-2, and M3M-3 respectively where $M$ is the maximum refinement level.
@@ 6. RAPTORCase Studies: Experimental Results and Insights for the four AMR Workloads
Hereafter, we describe our results from analyzing the Flash-X workloads with RAPTOR. We explore three truncation modes: (1) Global truncation, where numerical values are truncated uniformly across the entire domain, (2) Selective truncation with AMR, where truncation is applied only on levels coarser than Ml{M-l}, with $l$ controlling the cutoff depth in the AMR hierarchy, and (3) Selective truncation of a complete physics module.
@@ 6.1. Compressible hydrodynamics applications
Figures 7(a) and 7(b) show the results of the hydro experiments. As expected, the Sedov problem exhibits a very robust tolerance to lowered precision at M1M-1 as demonstrated in the second panel of Figure 7(a), which shows a much lower error for any size of mantissa compared to the error shown in the first panel. For mantissas with less than 28 bits, excluding the finest AMR blocks from truncation results in an error reduction of more than 7 orders-of-magnitude. For mantissas larger than 28 bits, the error remains constant and approaches the error measured for full truncation. This error floor appears as the simulation result drifts over time due to small inaccuracies in truncated floating-point operations. Excluding the second level of AMR blocks from truncation does not significantly improve the error of the simulation, as demonstrated by the unchanging error in Panel 3. Only when we exclude the finest three levels of AMR blocks from truncation do we see that the constant error floor drops by another half order of magnitude as shown in Panel 4.
@@
The colored bars in the background of the figure panels count the number of truncated and full-precision floating-point operations occurring for each simulation. In the fourth panel, the workload is comprised almost entirely of full-precision operations, with truncated operations making up less than $0.4\%$. In the third and second panel, the truncated operations make up progressively larger portions of the workload as the Hydro module is truncated for more blocks in the domain. Finally, in the first panel, the truncated operations make up more than 80% of the floating-point workload. Additionally, for mantissas smaller than 10 bits, the bars show irregularities, with the overall flop counts fluctuating. This occurs because the low-precision operations are influencing the decisions of the AMR algorithm, which is responsible for dynamically refining the mesh where needed. To be clear, it is not the algorithm itself which is working with truncated precision. Rather, the algorithm notices imprecise blocks and decides to refine them in order to maintain numerical accuracy and convergence in the simulation.
@@
Figure 7(b) shows the error plot for the Sod problem. Comparing the first two panels, we observe that excluding the finest AMR blocks from truncation results in an improvement in the overall error for mantissa sizes smaller than 18 bits. This improvement results in an almost order-of-magnitude difference in error between the first and second panel for simulations using 4 bits of mantissa. Panel 3 shows that excluding another level of AMR blocks from truncation results in an error improvement across the board, with the difference starting at about half an order-of-magnitude for large mantissas with 48 bits and increasing to almost an order-of-magnitude for small mantissas with 10 bits or fewer.
@@
Excluding refined blocks from truncation improves the error more for the Sedov problem than it does for the Sod problem, validating Hypothesis 1. For Sedov, running just the finest blocks at full precision results in an improvement of 7 orders-of-magnitude compared to truncating all blocks. For Sod, the same strategy decreases the error by an order-of-magnitude at most. The figure for Sod has one less panel, because eventually no leaf blocks remain at the M3M-3 level in the simulation, so no truncation occurs and therefore there is no error.
@@
In Figure 7(b), an anomalous behavior is visible for mantissas with 4-6 bits, where the errors are comparable to those obtained for mantissas with 20 bits. The reason for that anomaly is written in the logfile of the Flash-X run, i.e., it shows a higher number of leaf blocks (the blocks on which the solution evolves) which implies that many more blocks meet the refinement criterion with so few mantissa bits. Consequently, any advantage from lowering the precision further is more than offset by the increased number of refined blocks. It highlights the importance of a careful cost-benefit analysis of any performance-related tweaks in the application and showcases how RAPTOR can assist with such an analysis.
@@
With Cellular we turn our attention to the possibility of truncating an entire physics module. The EOS module is the obvious choice here because the ordinary differential equations (ODEs) in the Burn module are particularly stiff and sensitive to numerical perturbation. Hypothesis 2 is based on the assumption that the computation can work with reduced precision because it extrapolates from a table look-up. However, this hypothesis is proven to be incorrect because the Newton-Raphson algorithm, used to extrapolate from the table look-up values to actual conditions in the simulation, does not converge within the specified number of iterations when the mantissa is truncated to less than 42 bits. In an attempt to get around this problem, we decrease the tolerance for convergence and increase the permitted number of iterations. Yet, we fail to get convergence for any meaningful workload, providing us with a counterexample for our hypothesis. This experiment in particular highlights the usefulness of RAPTOR in numerical profiling because scientific intuition is not always reliable. Any direct modification of code to run such experiments with reduced precision would be time consuming and not as informative.